Assume that a random variable. X, follow a chi-square distribution with 23 degrees of freedom. Then

use this information to determine the following:

(a) P(14.85 < X<32.01).

(b) Constants and b such that Pa< X<b) = 0.95 and PX a) = 0.025. (c) The mean and variance of X.
1
Expert's answer
2021-09-02T09:01:21-0400

(a) "\nu=23"


"P(14.85<X<32.01)=P(X<32.01)"

"-P(X\leq14.85)=0.900064-0.100071"

"=0.799993"



(b) "\nu=23"


"P(X<a)=0.025"

"a=18.1373"

"P(a<X<b)=P(X<b)-P(X\leq a)"

"=P(X<b)-0.025=0.95"

"P(X<b)=0.975"

"b=38.075"

(c) "\nu=23"

The mean and variance of the chi-squared distribution are


"\mu=\nu=23"

"Var(X)=\sigma^2=23^2=529"